ToxicaDennisIce:

arbara used the distributive property to write an expression that is equivalent to 7 d + 63. Her work is shown below. The greatest common factor is 7 so 7 d + 63 = 7 (d + 63) What was Barbara’s error? A Barbara wrote (d + 63) instead of 63 d). B Barbara wrote (d + 63) instead of (d + 9). C Barbara wrote (d + 63) instead of (d + 56). D Barbara wrote (d + 63) instead of (10 d).

ToxicaDennisIce:

help?

jimthompson5910:

What happens when you distribute 7(d + 63)?

ToxicaDennisIce:

uh

Timmyspu:

So which one do you think is a wrong awnser

ToxicaDennisIce:

C

Timmyspu:

ok good job anything else

jimthompson5910:



ToxicaDennisIce:

sure

CocoCammron:

No i think its b

ToxicaDennisIce:

so C was the right answer

CocoCammron:

B-Barbara wrote (d+63) Instead of (d+9) Step-by-step explanation: Since the greatest common factor is 7 it simply mean 7 is divisor of the expression 7d +63 and not just a divisor of 7d. =(7d+63)/7 = (d + 9) Expression 7(d+63) = (7*d + 7*63) = 7d + 441 which is not equal to 7d +63 So Barbara should have writen 7d +63 = 7(d+9) instead.

ToxicaDennisIce:

ok ok

CocoCammron:

Can you give me a medal

ToxicaDennisIce:

ok ok the answer was

CocoCammron:

??

ToxicaDennisIce:

medal?

ToxicaDennisIce:

for u

CocoCammron:

yeah

ToxicaDennisIce:

no

ToxicaDennisIce:

naw im jp

Timmyspu:

that is agenst the rules to ask for medals

ToxicaDennisIce:

it is

CocoCammron:

OOF

Timmyspu:

i could report you for that

ToxicaDennisIce:

look what my screen said

CocoCammron:

Ok sorry

ToxicaDennisIce:

CocoCammron:

Wow I think it was @Timmyspu who did it

ToxicaDennisIce:

He's not a moderator

Timmyspu:

no i did not do that

CocoCammron:

Who did then?

ToxicaDennisIce:

idk???